OpenStudy (anonymous):

find the real numbers x and y to make the equation true 8x+8yi=16+24i

OpenStudy (anonymous):

8x = 16 8y = 24

OpenStudy (whpalmer4):

This is a simple "make them match" problem. \[8x+8yi = 16 + 24i\]Clearly, to make the \(i\) terms match, the following must be true:\[8yi = 24i\]Solve for \(y\) To make the other two terms match, the following must be true:\[8x=16\]Solve for \(x\)

OpenStudy (cwrw238):

the real and imaginary parts should be equal

OpenStudy (whpalmer4):

The two real parts should be equal, and the two imaginary parts should be equal, but they will not be equal to each other :-)

OpenStudy (anonymous):

also this one same thing 2x-7yi=-14+21i

OpenStudy (anonymous):